OpenStudy (anonymous):

Identify whether the series is a convergent or divergent geometric series and find the sum, if possible. (2 points)

OpenStudy (campbell_st):

find the common ratio. if \[-1 < r < 1\] then the series is convergent if its convergent you can use the infinite sum formula \[S_{\infty} = \frac{a}{1 - r}\] hope it helps

OpenStudy (anonymous):

Summation from 1 to infinity 16(5)^ \[\sum_{1}^{\infty}16(5)^{i-1}\]i-1

OpenStudy (campbell_st):

ok... well this will be divergent... look at the pattern of terms i = 1 16 i = 2 80 i = 3 400 etc... so since r > 1 the series is divergent